# Gravitational wave propagation in bigravity in the late universe

### v1

@@ 3.1 Regimes of physical interest
Figure 1: Plots of the ratio of the gravitational-wave luminosity distance to the electromagnetic luminosity distance in the ‘temporally unresolved’ regime, as a function of the redshift of the emitting source given by (3.17). The behavior changes significantly depending on the value of the mass.
Specifically, for values of the mass such that $|\xi|^{2}\leq 2\pi k/H$, the ratio of luminosity distances (3.17) is a monotonically increasing function of the redshift $z_{0}$ and, for large redshift values, it approaches the limit
$\frac{d_{L}^{\text{gw}}(z_{0})}{d_{L}^{\text{em}}(z_{0})}\to\frac{1}{\cos^{2}\theta}\left[1+\tan^{4}\theta+2\tan^{2}\theta\cos\left(\frac{|\xi|^{2}H}{2k}\right)\right]^{-1/2}~{}.$ (3.19)
For $|\xi|^{2}>2\pi k/H$, (3.17) transitions to an oscillatory behavior, subject to the bounds $1\leq d_{L}^{\text{gw}}(z_{0})/d_{L}^{\text{em}}(z_{0})\leq|\cos(2\theta)|^{-1}$. Interestingly, the upper bound is divergent when the mixing angle is $\theta=\frac{\pi}{4}$. For $\theta\neq\frac{\pi}{4}$, the total number of peaks and troughs is $\left\lceil\frac{H|\xi|^{2}}{2\pi k}-1\right\rceil$.
In turn, in the regime where the massless and massive components can be resolved at the detector (‘temporally resolved’ regime), the luminosity distance is determined only by the massless component, represented by the first term in the solution (3.15b). (As explained later in this section, the massive component represents a delayed ‘echo’ of the original signal.) The amplitude of the massless component is suppressed by a factor $\cos^{2}\theta$ compared to the corresponding solution in GR. Hence, the luminosity distance in the ‘temporally resolved’ regime reads
@@
$\begin{split}&h_{2}(\eta,x)=\Re\left[\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k}{2\pi}\left(\frac{\eta}{\eta_{0}}\right)\left(1+ik\eta_{0}\right)e^{i\big{(}\Phi_{\textbf{k}}(\eta)-\Phi_{\textbf{k}}(\eta_{0})+k(x-x_{0})\big{)}}Ne^{-\frac{\sigma^{2}}{2}(k-k_{0})^{2}}\right]\\ &\approx N\left(\frac{\eta}{\eta_{0}}\right)\Re\left[e^{i\big{(}\Phi_{k_{0}}(\eta)-\Phi_{k_{0}}(\eta_{0})+k_{0}(x-x_{0})\big{)}}\int\limits_{-\infty}^{+\infty}\frac{\mbox{d}k}{2\pi}\left(1+ik\eta_{0}\right)e^{i(k-k_{0})\left(\sqrt{\eta^{2}+\frac{|\xi|^{2}}{k_{0}^{2}}}-\sqrt{\eta_{0}^{2}+\frac{|\xi|^{2}}{k_{0}^{2}}}+x-x_{0}\right)}e^{-\frac{\sigma^{2}}{2}(k-k_{0})^{2}}\right]\\ &=\frac{N\,e^{-\frac{W^{2}}{2\sigma^{2}}}}{\sqrt{2\pi}\sigma}\left[\left(\left(\frac{\eta}{\eta_{0}}\right)-\frac{\eta W}{\sigma^{2}}\right)\cos(k_{0}\widetilde{W})-k_{0}\eta\sin(k_{0}\widetilde{W})\right]~{},\end{split}$ (3.24)
@@
Figure 2: The figure depicts the evolution of wave-packet solutions for the $h$ (blue curve) and $\tilde{h}$ (green curve) modes in the ‘propagating regime’ in the ‘temporally unresolved’ case, for initial data such that $\tilde{h}(\eta_{0},x)=0$ and only right-moving modes present, corresponding to the analytical solutions (3.22), with parameters $|\xi|=10$, $\theta=\pi/6$, $H=10^{-1}$, $k_{0}=20H$, $\sigma=10$, $x_{0}=-50$, $\eta_{0}=-60$, $N=H$. The plots above are equally spaced in conformal time, starting at $\eta=\eta_{0}$ and increasing in time steps $\Delta\eta=\sigma$. The dominant effect that characterizes the propagation in real space is an oscillation phenomenon, which stems from the nontrivial relation (mixing) between $h$ and $\tilde{h}$ with the mass eigenstates
@@
Figure 3: Evolution of the massless $h_{1}$ (continuous red curve) and massive $h_{2}$ (dashed yellow curve) components of the solution (3.22) in the ‘temporally unresolved’ case, with the same initial conditions, parameters, and time spacing used in Figure 2(f). The separation between the two components of the signal, due to the subluminal group velocity of the massive one, is particularly evident in the bottom panels.
The evolution of the wave packets is illustrated in Figures 2(f) and 3. The group velocity and the phase velocity of the $h_{2}$ mode are given, respectively, by [^3] These results crucially depend on the fact that we have not approximated the phase function (3.11) at the outset. In fact, using the approximations (3.14) to build the wave packets would lead to the incorrect prediction of a superluminal group velocity.
@@
We observe that ${\rm v}_{g}<1$ at all times, whereas the phase velocity ${\rm v}_{p}$ is superluminal. In both cases, deviations from the speed of light are small, since in this regime we are assuming $|\xi|\ll k_{0}|\eta|$, but they increase with time as $|\eta|$ decreases towards the future. We can use the above results to compute the difference in arrival time $\Delta T$ between the massless and massive components, considering a propagation distance $L$ between the source and an observer at redshift $z=0$ (corresponding to $\eta=-1/H$)
@@
$L\gg L_{\rm res.}\equiv\sigma\frac{2k_{0}^{2}}{m_{\rm FP}^{2}}~{}.$ (3.28)
@@
In summary, in the ‘temporally resolved’ regime, the delayed copies (or ‘echoes’) of the gravitational wave signal carried by the $h_{2}$ mode undergo two main effects: a distortion of their temporal profile due to the mass $m_{\rm FP}$, and amplitude suppression by a factor $\sin^{2}\theta$. The analysis of the propagation of wave packets given above also clarifies the different regimes of applicability of the luminosity distance formulae, (3.17) and (3.20). We remark that the derivation of Eq. (3.17) only holds in the ‘temporally unresolved’ regime, where the massless and massive signal components are overlapping. However, for high redshift sources, such that condition (3.28) holds and one enters the ‘temporally resolved’ regime, the massless component is observed first at the detector and as a separate signal component, in which case the luminosity distance is given by Eq. (3.20).
@@
which describes a nonpropagating field. In terms of the original metric perturbations, we obtain
[…]
@@
Propagation of wave packets
with $U=\eta-x-\eta_{0}+x_{0}$, and
The evolution of the $h$ mode in this example is illustrated in Figure 4.
@@ 4 Decoherence of bigravity oscillations?
Regarding the detection of tensor modes in the bigravity framework, only $h$ is directly relevant, as it is the only one that couples to matter. In previous sections, we showed that $h$ propagates as a superposition of mass eigenstates: one massless, $u$, and one massive, $v$. These modes obey different dispersion relations and, as a result, the corresponding signals in real space propagate with different group velocities. In turn, this implies that, for signals with finite duration (such as wave packets) and sufficiently large propagation trajectories, there comes a point where the two signals no longer overlap and, from that point on, propagate without displaying any interference patterns. This phenomenon has been previously referred to in the literature as ‘decoherence of bigravity oscillations’ [22, 23]. However, in the absence of matter fields, since the system is purely classical, there is nothing that could cause two initially coherent wave packets to decohere. The purpose of this section is to show that, even in the presence of incoherent matter sources coupled to gravity, wave packets of the mass eigenstates can preserve their initial coherence even long after they separate. Moreover, unlike neutrino oscillations, correlation functions of mass eigenstates decay algebraically, and therefore, there is no well-defined notion of a coherence length associated with bigravity oscillations, contrary to previous claims.
@@ 4.1 Wave packets propagating in an incoherent medium
where $\omega_{k}\coloneqq\sqrt{k^{2}+m_{\rm FP}^{2}}$ If we restrict our attention to positive-frequency solutions of the inhomogeneous system (4.2), the general solution can be written as
@@
where $\eta_{0}$ represents some initial time where we set initial conditions (for instance, corresponding to the emission event, as we did in the previous section). The Fourier components of the matter source are subject to the reality conditions $\pi_{\textbf{k}}=\pi_{-\textbf{k}}^{*}$. On the contrary, it is not necessary to impose reality conditions on the homogeneous solution, represented by the first terms in (4.3a)–(4.3b), as we will take the real part of the Fourier integral to ensure reality of the wave-packet solution. For simplicity, in the following we assume propagation only in the $x$ direction (therefore, k-dependent functions only have support on wave vectors pointing in the positive $x$ direction, that is, $\textbf{k}=k_{x}\hat{\textbf{x}}$). Taking the inverse Fourier transform, we obtain the corresponding wave-packet solutions in position space
@@
$v(\eta,x)$ $=\Re\left[\int_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi}\,B_{\textbf{k}}e^{-i\left(\omega_{k}(\eta-\eta_{0})-k_{x}x+\phi_{\textbf{k}}\right)}\right]-\frac{2\sin\theta}{M_{g}^{2}}\int_{\eta_{0}}^{\eta}\mbox{d}\eta^{\prime}\int_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi}\,\frac{\sin(\omega_{k}(\eta-\eta^{\prime}))}{\omega_{k}}e^{ikx}\pi_{\textbf{k}}(\eta^{\prime})~{}.$ (4.4b)
@@ 4.2 Correlation functions of mass eigenstates
In this section, we systematically compute the correlation functions of wave-packet solutions for the mass eigenstates. Given two modes, $V(\eta^{\prime},x^{\prime})$ and $\widetilde{V}(\eta,x)$, evaluated at spacetime points $(\eta,x)$ and $(\eta^{\prime},x^{\prime})$, respectively, their correlation function is defined as
@@ 5 Discussion
The behavior of the massive mode changes dramatically depending on the value of the mass $m_{\rm FP}$ relative to the wavenumber $k$, and we identify two limiting regimes of interest. The first regime, analyzed in Section 3.1.1, is defined by the condition that the wavenumber is large compared to the mass, that is, $m_{\rm FP}/H\ll|\eta|k$, whereby the massive graviton is freely propagating. The relevant asymptotics of the mode functions are given in Eq. (3.14). Since the group velocity of the massive mode is subluminal, and thus it propagates more slowly than the massless mode, depending on the distance traveled in position space, one can identify two different sub-regimes, depending on whether the massive and massless modes can be temporally resolved or not. In the ‘temporally unresolved’ sub-regime, the massive mode causes a modulation of the gravitational wave signal. In particular, we obtained the analytical formula (3.17) for the gravitational-wave luminosity distance versus the electromagnetic luminosity distance, $d_{L}^{\rm gw}/d_{L}^{\rm em}$, expressed as a function of the redshift of the emitting source. The behavior of this function varies significantly, depending on the value of the mass. Specifically, we identify a threshold value for the mass (3.18), which was missed in previous work [5]. For sub-threshold values $d_{L}^{\rm gw}/d_{L}^{\rm em}$ is a monotonically increasing function of redshift, whereas it transitions to an oscillatory behavior for above-threshold values, as shown in Figure 1. In both cases, the upper bound is finite and determined solely by the mixing angle (except for the special case $\theta=\frac{\pi}{4}$, where said upper bound is divergent). However, when the propagation distance is sufficiently large—as specified by the condition (3.28)—the massive and massless components separate, and one enters the ‘temporally resolved’ sub-regime. In this case, the massless and massive modes represent two distinct components of the signal. The amplitude of the massless mode is suppressed by a $\cos^{2}\theta$ factor compared to a GR solution subject to the same initial data, while the amplitude of the delayed ‘echo’ carried by the massive mode is suppressed by a $\sin^{2}\theta$ factor. Moreover, unlike the massless mode, whose profile is undistorted, the temporal profile of the gravitational wave ‘echo’ is broadened and distorted due to propagation. The gravitational-wave luminosity distance in this regime is determined solely by the massless component, and the value of $d_{L}^{\rm gw}/d_{L}^{\rm em}$ is a simple function of the mixing angle, as given in Eq. (3.20).
@@
We remark that the ‘temporally resolved’ and ‘temporally unresolved’ regimes have previously been referred to in the literature as the ‘coherent’ and ‘decoherent’ regimes, respectively, see Refs. [22, 23]. However, we find this terminology potentially misleading and thus deliberately avoid it in our work. As explicitly shown in Section 4, the massless and massive modes remain coherent even in the regime where they can be temporally resolved. We analytically computed the correlation and autocorrelation functions of the mass eigenstates, showing that they decay algebraically for large separations between spacetime events. Therefore, it is not possible to define a coherence length associated with bigravity oscillations. This shows the limitation of the analogy between the classical phenomenon of bigravity oscillations and neutrino oscillations (where the correlations are inherently quantum and decay exponentially as a function of the separation between wave packets of the mass eigenstates).
@@
This work can be extended in several directions. In order to extend our analysis to the propagation of gravitational waves emitted by higher-redshift sources, which were active before the late-time de Sitter era, we need to consider more general cosmological solutions on the viable branch of bimetric gravity. In this more general scenario, neither the ratio of the scale factors of the two metrics nor their relative lapse is constant. As a result, the system of equations (2.6) cannot be decoupled in terms of mass eigenstates at higher redshift. To address this, appropriate techniques based on multiple-scale analysis [31, 42] must be employed to account for the influence of the evolving cosmological background on the dynamics of tensor modes, which smoothly reduces to the de Sitter case analyzed in this paper. Another natural extension of this work would be to consider multigravity theories involving $N\geq 3$ conformally related background metrics [43, 44], with tensor perturbations corresponding to linear combinations of one massless and $N-1$ massive gravitons. In this context, it would be interesting to identify possible mass hierarchies, classify different dynamical regimes that may arise, and study their imprints on gravitational wave signals and the luminosity distance.

### v2

@@ 3.1 Regimes of physical interest
Figure 1: Plots of the ratio of the gravitational-wave luminosity distance to the electromagnetic luminosity distance in the ‘temporally unresolved’ regime, as a function of the redshift $z_{0}$ of the emitting source given by (3.17). The behavior changes significantly depending on the value of the mass.
Figure 2: The shaded area represents the region of parameter space $(\theta,m_{\mathrm{FP}})$ ruled out by the constraint (3.26), for $h=0.7$ . The left panel shows the excluded region for $\theta\in(0,\pi/2)$ in a linear scale, while the right panel shows a close-up for $\theta\in(10^{-40},1)$ in a logarithmic scale to better illustrate the behavior of the contour for small angles.
@@
[…]
@@
Figure 3: The figure depicts the evolution of wave-packet solutions for the $h$ (blue curve) and $\tilde{h}$ (green curve) modes in the ‘propagating regime’ in the ‘temporally unresolved’ case, for initial data such that $\tilde{h}(\eta_{0},x)=0$ and only right-moving modes present, corresponding to the analytical solutions (3.28), with parameters $|\xi|=10$, $\theta=\pi/6$, $H=10^{-1}$, $k_{0}=20H$, $\sigma=10$, $x_{0}=-50$, $\eta_{0}=-60$, $N=H$. The plots above are equally spaced in conformal time, starting at $\eta=\eta_{0}$ and increasing in time steps $\Delta\eta=\sigma$. The dominant effect that characterizes the propagation in real space is an oscillation phenomenon, which stems from the nontrivial relation (mixing) between $h$ and $\tilde{h}$ with the mass eigenstates.
@@
Figure 4: Evolution of the massless $h_{1}$ (continuous red curve) and massive $h_{2}$ (dashed yellow curve) components of the solution (3.28) in the ‘temporally unresolved’ case, with the same initial conditions, parameters, and time spacing used in Figure 3. The two modes evolve independently, and both experience the usual amplitude damping due to the expansion of the Universe. Moreover, the massive mode experiences dispersion. The separation between the two components of the signal, due to the subluminal group velocity of the massive one, is particularly evident in the bottom panels.
The evolution of the wave packets is illustrated in Figures 3 and 4. The group velocity and the phase velocity of the $h_{2}$ mode are given, respectively, by [^6] These results crucially depend on the fact that we have not approximated the phase function (3.11) at the outset. In fact, using the approximations (3.14) to build the wave packets would lead to the incorrect prediction of a superluminal group velocity.
@@
We observe that ${\rm v}_{g}<1$ at all times, whereas the phase velocity ${\rm v}_{p}$ is superluminal. In both cases, deviations from the speed of light are small, since in this regime we are assuming $|\xi|\ll k_{0}|\eta|$, but they increase with time as $|\eta|$ decreases towards the future.
We can use the above results to compute the difference in arrival time $\Delta T$ between the massless and massive components, considering a propagation distance $L$ between the source and an observer at redshift $z=0$ (corresponding to $\eta=-1/H$)
@@
$L\gg L_{\rm res.}\equiv\sigma\frac{2k_{0}^{2}}{m_{\rm FP}^{2}}\sim\left(\frac{\sigma}{0.1\,{\rm s}}\right)\left(\frac{k_{0}}{100\,{\rm Hz}}\right)^{2}\left(\frac{10^{-22}\,{\rm eV}}{m_{\rm FP}}\right)^{2}{\rm Gpc}~.$ (3.34)
@@
In summary, in the ‘temporally resolved’ regime, the delayed copies (or ‘echoes’) of the gravitational wave signal carried by the $h_{2}$ mode undergo two main effects: a distortion of their temporal profile due to the mass $m_{\rm FP}$, and amplitude suppression by a factor $\sin^{2}\theta$. The analysis of the propagation of wave packets given above also clarifies the different regimes of applicability of the luminosity distance formulas, (3.17) and (3.24). We remark that the derivation of Eq. (3.17) only holds in the ‘temporally unresolved’ regime, where the massless and massive signal components are overlapping. However, for high redshift sources, such that condition (3.34) holds and one enters the ‘temporally resolved’ regime, the massless component is observed first at the detector and as a separate signal component, in which case the luminosity distance is given by Eq. (3.24).
@@
which describes a nonpropagating field. Note, however, that we are not necessarily constrained to super-horizon scales in this regime: for values of $m_{\rm FP}$ much greater than the Higuchi lower bound, one may well have $|\eta|k\gg 1$, corresponding to sub-horizon scales whereby the massless mode freely oscillates. In terms of the original metric perturbations, we obtain
$\begin{split}h_{\textbf{k}}(z)&\approx A_{\textbf{k}}\cos\theta\,\left(\cos\left(\frac{k}{H}(1+z)+\varphi_{\textbf{k}}\right)+\frac{k}{H}(1+z)\sin\left(\frac{k}{H}(1+z)+\varphi_{\textbf{k}}\right)\right)\\ &\quad-\hbox{\set@color\hskip 2.77779pt\hskip-2.13194pt\hbox{\set@color$\hskip-1.29167pt\widetilde{B}$}\hskip-2.13194pt\hskip-2.77779pt\raisebox{-0.95827pt}{\hbox{\set@color$\smash{\widetilde{\phantom{B}}}$}}\hskip-2.77779pt\hskip 2.77779pt}_{\textbf{k}}\sin\theta\,(1+z)^{3/2}\cos\left(|\xi|+|\xi|\ln\left(\frac{k(1+z)}{2H|\xi|}\right)+\phi_{\textbf{k}}\right)~,\end{split}$ (3.36a)
$\begin{split}\tilde{h}_{\textbf{k}}(z)&\approx A_{\textbf{k}}\cos\theta\,\left(\cos\left(\frac{k}{H}(1+z)+\varphi_{\textbf{k}}\right)+\frac{k}{H}(1+z)\sin\left(\frac{k}{H}(1+z)+\varphi_{\textbf{k}}\right)\right)\\ &\quad+\hbox{\set@color\hskip 2.77779pt\hskip-2.13194pt\hbox{\set@color$\hskip-1.29167pt\widetilde{B}$}\hskip-2.13194pt\hskip-2.77779pt\raisebox{-0.95827pt}{\hbox{\set@color$\smash{\widetilde{\phantom{B}}}$}}\hskip-2.77779pt\hskip 2.77779pt}_{\textbf{k}}\frac{\cos^{2}\theta}{\sin\theta}(1+z)^{3/2}\cos\left(|\xi|+|\xi|\ln\left(\frac{k(1+z)}{2H|\xi|}\right)+\phi_{\textbf{k}}\right)~.\end{split}$ (3.36b)
@@
Propagation of wave packets
with $U=\eta-x-\eta_{0}+x_{0}$, and
Note that the solution for $h_{1}$ in Eq. (3.39) coincides with that obtained in Eq. (3.29) for the propagating regime. This is a direct consequence of the fact that the solution for the massless mode $u$ is exact and therefore independent of the regime under consideration. By contrast, the massive mode $v$ has different asymptotics in each regime, and hence only $h_{2}$ is regime-dependent. The evolution of the $h$ mode in this example is illustrated in Figure 5.
@@ 4 Decoherence of bigravity oscillations?
Regarding the detection of tensor modes in the bigravity framework, only $h$ is directly relevant, as it is the only one that couples to matter. In previous sections, we showed that $h$ propagates as a superposition of mass eigenstates: one massless, $u$, and one massive, $v$. These modes obey different dispersion relations and, as a result, the corresponding signals in real space propagate with different group velocities. In turn, this implies that, for signals with finite duration (such as wave packets) and sufficiently large propagation trajectories, there comes a point where the two signals no longer overlap and, from that point on, propagate without displaying any interference patterns. This phenomenon has been previously referred to in the literature as ‘decoherence of bigravity oscillations’ [29, 30]. However, in the absence of matter fields, since the system is purely classical, there is nothing that could cause two initially coherent wave packets to decohere. [^8] Here we adopt the standard definition of spatiotemporal coherence for classical radiation, as used also in optics [54]. The purpose of this section is to show that, even in the presence of incoherent matter sources coupled to gravity, wave packets of the mass eigenstates can preserve their initial coherence even long after they separate. Moreover, unlike neutrino oscillations, correlation functions of mass eigenstates decay algebraically, and therefore, there is no well-defined notion of a coherence length associated with bigravity oscillations, contrary to previous claims.
@@ 4.1 Wave packets propagating in an incoherent medium
where $\omega_{k}\coloneqq\sqrt{k^{2}+m_{\rm FP}^{2}}$ . If we restrict our attention to positive-frequency solutions of the inhomogeneous system (4.2), the general solution can be written as
@@
where $\eta_{0}$ represents some initial time where we set initial conditions (for instance, corresponding to the emission event, as we did in the previous section). The Fourier components of the matter source are subject to the reality conditions $\pi_{\textbf{k}}=\pi_{-\textbf{k}}^{*}$. On the contrary, it is not necessary to impose reality conditions on the homogeneous solution, represented by the first terms in (4.3a)–(4.3b), as we will take the real part of the Fourier integral to ensure reality of the wave-packet solution. For simplicity, in the following we assume propagation only in the $x$ direction, that is $\textbf{k}=k_{x}\hat{\textbf{x}}$. Taking the inverse Fourier transform, we obtain the corresponding wave-packet solutions in position space
@@
$v(\eta,x)$ $=\Re\left[\int_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi}\,B_{\textbf{k}}e^{-i\left(\omega_{k}(\eta-\eta_{0})-k_{x}x+\phi_{\textbf{k}}\right)}\right]-\frac{2\sin\theta}{M_{g}^{2}}\int_{\eta_{0}}^{\eta}\mbox{d}\eta^{\prime}\int_{-\infty}^{+\infty}\frac{\mbox{d}k_{x}}{2\pi}\,\frac{\sin(\omega_{k}(\eta-\eta^{\prime}))}{\omega_{k}}e^{ik_{x}x}\pi_{\textbf{k}}(\eta^{\prime})~.$ (4.4b)
@@ 4.2 Correlation functions of mass eigenstates
In this section, we systematically compute the correlation functions of wave-packet solutions for the mass eigenstates. In order to assess whether decoherence takes place, it is crucial to investigate the decay properties of such correlation functions. Given two modes, $V(\eta^{\prime},x^{\prime})$ and $\widetilde{V}(\eta,x)$, evaluated at spacetime points $(\eta,x)$ and $(\eta^{\prime},x^{\prime})$, respectively, their correlation function is defined as
@@
As shown below, the correlation functions between mass eigenstates always decay algebraically in the limit of large spatiotemporal separations. This shows that the system does not undergo decoherence, which is associated with an exponential decay of the correlation functions.
@@ 5 Discussion
The behavior of the massive mode changes dramatically depending on the value of the mass $m_{\rm FP}$ relative to the wavenumber $k$, and we identify two limiting regimes of interest. The first regime, analyzed in Section 3.1.1, is defined by the condition that the wavenumber is large compared to the mass, that is, $m_{\rm FP}/H\ll|\eta|k$, whereby the massive graviton is freely propagating. The relevant asymptotics of the mode functions are given in Eq. (3.14). Since the group velocity of the massive mode is subluminal, and thus it propagates more slowly than the massless mode, depending on the distance traveled in position space, one can identify two different sub-regimes, depending on whether the massive and massless modes can be temporally resolved or not. In the ‘temporally unresolved’ sub-regime, the massive mode causes a modulation of the gravitational wave signal. In particular, we obtained the analytical formula (3.17) for the gravitational-wave luminosity distance versus the electromagnetic luminosity distance, $d_{L}^{\rm gw}/d_{L}^{\rm em}$, expressed as a function of the redshift of the emitting source. The behavior of this function varies significantly, depending on the value of the mass. Specifically, we identify a threshold value for the mass (3.18), which was missed in previous work [5]. For sub-threshold values $d_{L}^{\rm gw}/d_{L}^{\rm em}$ is a monotonically increasing function of redshift, whereas it transitions to an oscillatory behavior for above-threshold values, as shown in Figure 1. In both cases, the upper bound is finite and determined solely by the mixing angle (except for the special case $\theta=\frac{\pi}{4}$, where said upper bound is divergent). However, when the propagation distance is sufficiently large—as specified by the condition (3.34)—the massive and massless components separate, and one enters the ‘temporally resolved’ sub-regime. In this case, the massless and massive modes represent two distinct components of the signal. The amplitude of the massless mode is suppressed by a $\cos^{2}\theta$ factor compared to a GR solution subject to the same initial data, while the amplitude of the delayed ‘echo’ carried by the massive mode is suppressed by a $\sin^{2}\theta$ factor. Moreover, unlike the massless mode, whose profile is undistorted, the temporal profile of the gravitational wave ‘echo’ is broadened and distorted due to propagation. The gravitational-wave luminosity distance in this regime is determined solely by the massless component, and the value of $d_{L}^{\rm gw}/d_{L}^{\rm em}$ is a simple function of the mixing angle, as given in Eq. (3.24). We remark that the phenomenological parametrization of the luminosity distance proposed in Ref. [51] does not apply to bigravity. We also showed that the independent measurements of the luminosity distance obtained with multimessenger event GW170817 can be used to place a constraint (3.26) in the parameter space of the theory $(\theta,m_{\rm FP})$, which still allows for significant deviations from GR at high redshift. The corresponding exclusion plot is shown in Figure 2.
@@
We remark that the ‘temporally resolved’ and ‘temporally unresolved’ regimes have previously been referred to in the literature as the ‘coherent’ and ‘decoherent’ regimes, respectively, see Refs. [29, 30]. However, we find this terminology potentially misleading and thus deliberately avoid it in our work. As explicitly shown in Section 4, the massless and massive modes remain coherent even in the regime where they can be temporally resolved. We analytically computed the correlation and autocorrelation functions of the mass eigenstates, showing that they decay algebraically for large separations between spacetime events. Therefore, it is not possible to define a coherence length associated with bigravity oscillations. This shows the limitation of the analogy between the classical phenomenon of bigravity oscillations and neutrino oscillations (where the correlations are inherently quantum and decay exponentially as a function of the separation between wave packets of the mass eigenstates). This effect is in principle a testable prediction of the theory: if both temporally-resolved components are observed at the detector, the spatiotemporal coherence of the gravitational wave signal may be probed through the autocorrelation function of the gravitational-wave strain time series.
@@
This work can be extended in several directions. In order to extend our analysis to the propagation of gravitational waves emitted by higher-redshift sources, which were active before the late-time de Sitter era, we need to consider more general cosmological solutions on the viable branch of bimetric gravity. In this more general scenario, neither the ratio of the scale factors of the two metrics nor their relative lapse is constant. As a result, the system of equations (2.6) cannot be decoupled in terms of mass eigenstates at higher redshift. To address this, appropriate techniques based on multiple-scale analysis [44, 59] must be employed to account for the influence of the evolving cosmological background on the dynamics of tensor modes, which smoothly reduces to the de Sitter case analyzed in this paper. Another natural extension of this work would be to consider multigravity theories involving $N\geq 3$ conformally related background metrics [60, 61], with tensor perturbations corresponding to linear combinations of one massless and N1N-1 massive gravitons. In this context, it would be interesting to identify possible mass hierarchies, classify different dynamical regimes that may arise, and study their imprints on gravitational wave signals and the luminosity distance.